Please help and explain all steps9 Marks [5 0 0 1 8. Let A= 10 3 [0 0 -2] (a) Find all eigenvalues of A and their corresponding eigenvectors. (b) Is A diagona
Answer #1

A - 0 0 7 5 LO 0 -2) A DI] = 15-1 0 0 1 .15- 0 0 1 11 - 3 To 0 -2-1) 14-051 =(5-1): -2-01 = (5-4) (-1|-2-)] - (5-) 1 (2+1) ThRis BRA TiO 3 O 0 -2 Rg- Ri R₂ rio 01 0 0 3 0-2 232 OOOWO -R2 + R3 1 0 1 3 Schue lo : 186 Sohee I looirien LoOo URO, BRO =0 UTi 0 0 1 - loool Ra R2 ſ! 0 0 1 1 3121 14100 11001 To 3/2 Ilu, Lo o o J и у 0,50 02 + 2 Uz =0 Ug= -303 Let Uzak l2=-2k 666 %Sotut RIER2 Il-50 Rgb R3 80he si -5 orien re,- 50g=0= 0,= 502 13-0 Let Ug=k 6,=5k kl leis k 9 OJ LJ LOJ (8) 17818) hvala eige